# Morse-Bott inequalities, Topology Change and Cobordisms to Nothing

### v1

@@ 6 Conclusions and Outlook
Acknowledgments: We are very grateful to Matilda Delgado, Miguel Montero and Irene Valenzuela for their invaluable contributions through extensive discussions and comments on the manuscript. We would like to also thank Alberto Castellano, Markus Dierigl, Bjoern Friedrich, Arthur Hebecker, Jesús Huertas, Andriana Makridou, Luca Melotti, Jacob McNamara, Jakob Moritz, Eduardo Velasco and Matteo Zatti for illuminating discussions and comments along the past two years. I.R. wishes to acknowledge the hospitality of the Depart of Theoretical Physics at CERN and the Department of Physics of Harvard University during different stages of this work. I.R. acknowledges the support of the Spanish FPI grant No. PRE2020-094163 and the ERC Starting Grant QGuide101042568 - StG 2021, as well as the Spanish Agencia Estatal de Investigación through the grant “IFT Centro de Excelencia Severo Ochoa” CEX2020-001007-S and the grant PID2021-123017NB-I00, funded by MCIN/AEI/10.13039/ 501100011033 and by ERDF “A way of making Europe”.
@@ Appendix A Basics of Compact Riemann Surfaces
In this appendix we review the basic results about Riemann surfaces that are used in section 4.2. For our purposes, a (compact) genus-$g$ Riemann surface $\Sigma_{g}$ can be seen as the connected sum of $g$ 2-torus $\mathbb{T}^{2}=\mathbb{S}^{1}\times\mathbb{S}^{1}$, $\Sigma_{g}:=\#^{g}\mathbb{T}^{2}$, with $\Sigma_{0}=\#^{0}\mathbb{T}^{2}\simeq\mathbb{S}^{2}$. Regardless of its genus, every compact Riemann surface $\Sigma$ admits a conformal Riemann metric Jost:2006 , i.e., $\mathrm{d}s_{\Sigma}^{2}=\lambda(z)^{2}\mathrm{d}z\mathrm{d}\bar{z}$, with $\lambda\in\mathcal{C}^{2}(\Sigma)$ and $z=x+iy$. A simple computation yields that the Ricci scalar is given by $\mathcal{R}_{\Sigma}=-2\Delta\log\lambda$, with $\Delta=\lambda^{-2}(\partial_{x}^{2}+\partial_{y}^{2})$. Given $\Gamma$ a discrete subgroup of
@@
acting discontinuously and without fixed points on the upper half plan =$\mathbb{H}=\{z=x+iy\in\mathbb{C}:\,y>0\}$, then $\Sigma=\mathbb{H}/\Gamma$ is a compact Riemann surface. Then, for an arbitrary $z_{0}\in\mathbb{H}$, the set
@@
where $d$ is the geodesic distance associated to the natural hyperbolic metric $\mathrm{d}s_{\mathbb{H}}^{2}=\frac{\mathrm{d}z\mathrm{d}\bar{z}}{2{\rm Im}(z)^{2}}$, is a convex polygon with finitely many sides, such that all the vertices are equivalent. It can then be shown Jost:2006 that sides of $F$ can be paired $a\equiv a^{\prime}$ uniquely, in such a way that the sides can be ordered as either $a_{0}a_{0}^{\prime}$ or $a_{1}b_{1}a^{\prime}_{1}b^{\prime}_{1}\dots a_{g}b_{g}a^{\prime}_{g}b^{\prime}_{g}$ in such a way that $F$ has $2$ or $4g$ sides. The fundamental group of $\Sigma$ (which is obtained from $F$ after sides are identified) is then
@@
For $g\geq 1$ we can take dual canonical homology and cohomology basis $\{a_{i},b_{i}\}_{i=1}^{g}$ and $\{\alpha^{i},\beta^{i}\}_{i=1}^{g}$ such that we have the following intersections and integrals Saltman:2004jh :
@@
On the other hand, from the Riemann-Roch theorem Jost:2006 ; Blumenhagen:2013fgp , one has that
#complex moduli parameters#conformal Killing $\#\text{complex moduli parameters}-\#\text{conformal Killing vectors}=-\frac{3}{2}\chi(\Sigma_{g})=3(g-1)\;,$ (A.8)
@@
  • For $\Sigma_{1}=\mathbb{T}^{2}$, we have a flat metric $\mathrm{d}s^{2}=\mathrm{d}z\mathrm{d}\bar{z}$. The translations along the torus are given by $U(1)\times U(1)$, with a single complex generator, resulting in one conformal Killing vector and one complex moduli parameter, $\tau=\tau_{1}+i\tau_{2}\in\mathbb{H}$. As one considers the $SL(2,\mathbb{Z})$ redundancies, the actual fundamental domain is smaller and corresponding with the famous =$\mathcal{F}=\left\{-\frac{1}{2}\leq\tau_{1}\leq 0\,,\,|\tau|^{2}\geq 1\right\}\cup\left\{0<\tau\leq\frac{1}{2}\,,\,|\tau|^{2}>1\right\}$.
[…]
@@
After the previous machinery has been introduced, we will now comment on two specific results from the literature that are used in section 4.2 and Appendix B . First of all, from the Uniformization theorem of Schwarz and Klein (see section 4.4 from Jost:2006 for proof in modern language), every compact Riemann surface admits constant curvature metric. This is indeed useful, as from Gauss-Bonnet theorem we have
@@
feature a Ricci flow (see Perelman:2006un ; Velazquez:2022eco and references therein)
@@
Finally, we will introduce some machinery in order that we will use for flux potentials. Given a 1-form $F\in\Omega^{1}(\Sigma_{g})$, we can expand it in both the dual and holomorphic basis as Saltman:2004jh
@@
In this appendix we give some basic overview of handle decay in Riemann surfaces, i.e., $\Sigma_{g}\to\Sigma_{g-1}$ via tachyon condensation. For more details, see Saltman:2004jh ; Adams:2005rb . We will work in the weak coupling and large volume limit ($g_{s}\ll 1$ and $\mathcal{V}\gg\ell_{s}^{2}$), so that string interactions and corrections can be safely ignored. As we will discuss later this section, this is the regime towards which a compactification on $\Sigma_{g}$ with $g\geq 2$ naturally evolves. From (4.9), $\Omega^{\rm Spin}_{1}=\langle\mathbb{S}_{p}^{1}\rangle\simeq\mathbb{Z}_{2}$, so 1-cycles need anti-periodic boundary conditions in order to shrink them to nothing (otherwise a spin defect would need to be included [^24] From McNamara:2019rup , in type the IIA $\mathbb{S}^{1}_{p}$ class in $\Omega_{1}^{\rm Spin}$ is killed through a $\mathbb{R}/\mathbb{Z}_{2}\times\mathbb{S}^{1}_{p}$ defect consisting in an O8-plane wrapped along the circle (in the bulk, not in the boundary to nothing). For type IIB, the T-dual of the defect has a (×$(\mathbb{R}\times\mathbb{S}^{1}_{p})/\mathbb{Z}_{2}$ geometry, with two O7-planes located at each end of the interval. These introduced defects are stringy by nature but, being located in the bulk of the bordism, should still allow for a smooth bordism to nothing after them.).
As shown in Appendix A, Ricci flow results in $\Sigma_{g}$ evolving towards having constant-curvature. For large volume, Gauss-Bonnet theorem implies $\alpha^{\prime}\mathcal{R}_{n}=2\pi(1-g)\frac{\alpha^{\prime}}{\mathcal{V}}\ll 1$, so that we can consider the topology change locus neighborhood to be approximately flat, locally looking like a SUSY-breaking Scherk-Schwarz circle Scherk:1978ta ; Scherk:1979zr ; Kounnas:1989dk of radius $\mathsf{R}$ with anti-periodic boundary conditions times a small interval, parameterized by $(\theta,r)\in\mathbb{S}^{1}\times(r_{0}-\epsilon,r_{0}+\epsilon)$. The world-sheet energy of the $n$-th twisted sector is given by
@@
so that for $\mathsf{R}<\sqrt{2\alpha^{\prime}}$ the spectrum will have tachyonic states, resulting in a instability in the internal geometry. As shown in footnote 18, a flux potential automatically stabilizes the complex structure moduli at finite values, so $\mathsf{R}<\sqrt{2\alpha^{\prime}}$ is energetically obstructed (though as seen in section 4.2, topology changes are energetically favored, at least in the studied settings). As we will argue now, this tachyon condensation results in a barrier that prevents propagation trough the $\mathbb{S}^{1}\times(r_{0}-\epsilon,r_{0}+\epsilon)$ tube, effectively changing the topology of our internal geometry.
@@
$S\supseteq\int\mathrm{d}^{2}\sigma\mathrm{d}\vartheta^{+}\mathrm{d}\vartheta^{-}T(X)\;,\quad\text{where}\quad T(X)=e^{k_{0}X^{0}}\hat{T}(R)\cos(\omega\tilde{\Theta})\,,$ (B.2)
with $X^{0}$, $R$ and $\tilde{\Theta}$ are the superspace coordinates associated to time $t$, radial direction $r$ and the T-dual $\bar{\theta}$ to $\theta$ Adams:2005rb , with $\hat{T}(R)$ complex as we have both winding and anti-winding modes in $T(X)$. It is further assumed that $\hat{T}(R)$ is localized at $r=r_{0}$, where the tube radius is the smallest, with exponentially decaying profile in the $r$ direction away from there. This results in a (classical) bosonic potential in the world-sheet action with the form
@@
Now, for $\omega^{2}\gg k_{r}^{2}>k_{0}^{2}$, this results in a classical barrier for all values of $\tilde{\theta}$ and $r$ over which $\hat{T}$ has support, precisely when $\partial_{r}\hat{T}$ becomes significant. This prevents the different world-sheet modes from propagating along the tube, effectively disconnecting the two sides of the handle and changing the topology of the internal manifold. This qualitative behavior remains after quantum effects are considered, see sections 2 and 3 from Adams:2005rb .

### v2

@@ 6 Conclusions and Outlook
Acknowledgments: We are very grateful to Matilda Delgado, Miguel Montero and Irene Valenzuela for their invaluable contributions through extensive discussions and comments on the manuscript. We would like to also thank Alberto Castellano, Markus Dierigl, Björn Friedrich, Arthur Hebecker, Camilo las Heras, Jesús Huertas, Andriana Makridou, Luca Melotti, Jacob McNamara, Jakob Moritz, Michelangelo Tartaglia, Eduardo Velasco and Matteo Zatti for illuminating discussions and comments along the past two years. I.R. wishes to acknowledge the hospitality of the Department of Theoretical Physics at CERN and the Department of Physics of Harvard University during different stages of this work. I.R. acknowledges the support of the Spanish FPI grant No. PRE2020-094163 and the ERC Starting Grant QGuide101042568 - StG 2021, as well as the Spanish Agencia Estatal de Investigación through the grant “IFT Centro de Excelencia Severo Ochoa” CEX2020-001007-S and the grant PID2021-123017NB-I00, funded by MCIN/AEI/10.13039/ 501100011033 and by ERDF “A way of making Europe”.
@@ Appendix A Basics of Compact Riemann Surfaces
In this appendix we review the basic results about Riemann surfaces that are used in section 4.2. For our purposes, a (compact) genus-$g$ Riemann surface $\Sigma_{g}$ can be seen as the connected sum of $g$ 2-torus $\mathbb{T}^{2}=\mathbb{S}^{1}\times\mathbb{S}^{1}$, $\Sigma_{g}:=\#^{g}\mathbb{T}^{2}$, with $\Sigma_{0}=\#^{0}\mathbb{T}^{2}\simeq\mathbb{S}^{2}$. Regardless of its genus, every compact Riemann surface $\Sigma$ admits a conformal Riemann metric Jost (2006), i.e., $\mathrm{d}s_{\Sigma}^{2}=\lambda(z)^{2}\mathrm{d}z\mathrm{d}\bar{z}$, with $\lambda\in\mathcal{C}^{2}(\Sigma)$ and $z=x+iy$. A simple computation yields that the Ricci scalar is given by $\mathcal{R}_{\Sigma}=-2\Delta\log\lambda$, with $\Delta=\lambda^{-2}(\partial_{x}^{2}+\partial_{y}^{2})$. Given $\Gamma$ a discrete subgroup of
@@
acting discontinuously and without fixed points on the upper half plan $\mathbb{H}=\{z=x+iy\in\mathbb{C}:\,y>0\}$, then $\Sigma=\mathbb{H}/\Gamma$ is a compact Riemann surface. Then, for an arbitrary $z_{0}\in\mathbb{H}$, the set
@@
where $d$ is the geodesic distance associated to the natural hyperbolic metric $\mathrm{d}s_{\mathbb{H}}^{2}=\frac{\mathrm{d}z\mathrm{d}\bar{z}}{2{\rm Im}(z)^{2}}$, is a convex polygon with finitely many sides, such that all the vertices are equivalent. It can then be shown Jost (2006) that edges of $F$ can be paired $a\equiv a^{\prime}$ uniquely, in such a way that the sides can be ordered as either $a_{0}a_{0}^{\prime}$ or $a_{1}b_{1}a^{\prime}_{1}b^{\prime}_{1}\dots a_{g}b_{g}a^{\prime}_{g}b^{\prime}_{g}$ in such a way that $F$ has $2$ or $4g$ sides. The fundamental group of $\Sigma$ (which is obtained from $F$ after sides are identified) is then
@@
For $g\geq 1$ we can take dual canonical homology and cohomology basis $\{a_{i},b_{i}\}_{i=1}^{g}$ and $\{\alpha^{i},\beta^{i}\}_{i=1}^{g}$ such that we have the following intersections and integrals Saltman and Silverstein (2006):
@@
On the other hand, from the Riemann-Roch theorem Jost (2006); Blumenhagen et al. (2013), one has that
$\text{complex moduli parameters}-\#\text{conformal Killing vectors}=-\tfrac{3}{2}\chi(\Sigma_{g})=3(g-1)\;,$ (A.8)
@@
  • For $\Sigma_{1}=\mathbb{T}^{2}$, we have a flat metric $\mathrm{d}s^{2}=\mathrm{d}z\mathrm{d}\bar{z}$. The translations along the torus are given by $U(1)\times U(1)$, with a single complex generator, resulting in one conformal Killing vector and one complex moduli parameter, $\tau=\tau_{1}+i\tau_{2}\in\mathbb{H}$. As one considers the $SL(2,\mathbb{Z})$ redundancies, the actual fundamental domain is smaller and corresponding with the famous $\mathcal{F}=\left\{-\frac{1}{2}\leq\tau_{1}\leq 0\,,\,|\tau|^{2}\geq 1\right\}\cup\left\{0<\tau\leq\frac{1}{2}\,,\,|\tau|^{2}>1\right\}$.
  • For $g\geq 2$ there are no conformal isometries, with the number of complex moduli parameters being 3g33g-3. While for g=2, 3g=2,\,3 this is the same as $\frac{1}{2}g(g+1)$, for $g\geq 4$ the Teichmüller space $\mathcal{T}_{g}$ is embedded in $\mathbb{H}^{g}$ in a non trivial way Jost (2006), so that no all choices of $\tau\in\mathbb{H}^{g}$ actually correspond to a complex structure in $\Sigma_{g}$. In any case, given $\Sigma_{g}$ and its homology classes, $\tau$ is uniquely determined. In general, we will have the same problem of defining the fundamental region $\mathcal{F}_{g}$ from $\mathcal{T}_{g}$, with Blumenhagen et al. (2013)
@@
After the previous machinery has been introduced, we will now comment on two specific results from the literature that are used in section 4.2 and Appendix B . First of all, from the Uniformization theorem of Schwarz and Klein (see section 4.4 from Jost (2006) for proof in modern language), every compact Riemann surface admits constant curvature metric. This is indeed useful, as from Gauss-Bonnet theorem we have
@@
feature a Ricci flow (see Perelman (2006); Velázquez et al. (2023) and references therein)
@@
Finally, we will introduce some machinery in order that we will use for flux potentials. Given a 1-form $F\in\Omega^{1}(\Sigma_{g})$, we can expand it in both the dual and holomorphic basis as Saltman and Silverstein (2006)
@@
In this appendix we give some basic overview of handle decay in Riemann surfaces, i.e., $\Sigma_{g}\to\Sigma_{g-1}$ via tachyon condensation. For more details, see Saltman and Silverstein (2006); Adams et al. (2005). We will work in the weak coupling and large volume limit ($g_{s}\ll 1$ and $\mathcal{V}\gg\ell_{s}^{2}$), so that string interactions and corrections can be safely ignored. As we will discuss later this section, this is the regime towards which a compactification on $\Sigma_{g}$ with $g\geq 2$ naturally evolves. From (4.9), $\Omega^{\rm Spin}_{1}=\langle\mathbb{S}_{p}^{1}\rangle\simeq\mathbb{Z}_{2}$, so 1-cycles need anti-periodic boundary conditions in order to shrink them to nothing (otherwise a spin defect would need to be included [^26] From McNamara and Vafa (2019), in type the IIA $\mathbb{S}^{1}_{p}$ class in $\Omega_{1}^{\rm Spin}$ is killed through a $\mathbb{R}/\mathbb{Z}_{2}\times\mathbb{S}^{1}_{p}$ defect consisting in an O8-plane wrapped along the circle (in the bulk, not in the boundary to nothing). For type IIB, the T-dual of the defect has a $(\mathbb{R}\times\mathbb{S}^{1}_{p})/\mathbb{Z}_{2}$ geometry, with two O7-planes located at each end of the interval. These introduced defects are stringy by nature but, being located in the bulk of the bordism, should still allow for a smooth bordism to nothing after them.).
As shown in Appendix A, Ricci flow results in $\Sigma_{g}$ evolving towards having constant-curvature. For large volume, Gauss-Bonnet theorem implies $\alpha^{\prime}\mathcal{R}_{n}=2\pi(1-g)\frac{\alpha^{\prime}}{\mathcal{V}}\ll 1$, so that we can consider the topology change locus neighborhood to be approximately flat, locally looking like a SUSY-breaking Scherk-Schwarz circle Scherk and Schwarz (1979b, a); Kounnas and Rostand (1990) of radius $\mathsf{R}$ with anti-periodic boundary conditions times a small interval, parameterized by $(\theta,r)\in\mathbb{S}^{1}\times(r_{0}-\epsilon,r_{0}+\epsilon)$. The world-sheet energy of the $n$-th twisted sector is given by
@@
so that for $\mathsf{R}<\sqrt{2\alpha^{\prime}}$ the spectrum will have tachyonic states, resulting in a instability in the internal geometry. As shown in footnote 20, a flux potential automatically stabilizes the complex structure moduli at finite values, so $\mathsf{R}<\sqrt{2\alpha^{\prime}}$ is energetically obstructed (though as seen in section 4.2, topology changes are energetically favored, at least in the studied settings). As we will argue now, this tachyon condensation results in a barrier that prevents propagation trough the $\mathbb{S}^{1}\times(r_{0}-\epsilon,r_{0}+\epsilon)$ tube, effectively changing the topology of our internal geometry.
@@
[…]
with $X^{0}$, $R$ and $\tilde{\Theta}$ are the superspace coordinates associated to time $t$, radial direction $r$ and the T-dual $\bar{\theta}$ to $\theta$ Adams et al. (2005), with $\hat{T}(R)$ complex as we have both winding and anti-winding modes in $T(X)$. It is further assumed that $\hat{T}(R)$ is localized at $r=r_{0}$, where the tube radius is the smallest, with exponentially decaying profile in the $r$ direction away from there. This results in a (classical) bosonic potential in the world-sheet action with the form
@@
Now, for $\omega^{2}\gg k_{r}^{2}>k_{0}^{2}$, this results in a classical barrier for all values of $\tilde{\theta}$ and $r$ over which $\hat{T}$ has support, precisely when $\partial_{r}\hat{T}$ becomes significant. This prevents the different world-sheet modes from propagating along the tube, effectively disconnecting the two sides of the handle and changing the topology of the internal manifold. This qualitative behavior remains after quantum effects are considered, see sections 2 and 3 from Adams et al. (2005).